OpenStudy (anonymous):

A line has a slope of 7 and contains the point (5, −9). Which equations represent the line? Choose ALL answers that are correct. A.−7x + y = −44 B.7x – y = 44 C.(y − 9) = 7(x + 5) D.(y + 9) = 7(x − 5) I'm pretty sure B. is correct and A. isn't, but I'm not sure about C. and D. .. help ?

OpenStudy (anonymous):

hey so how have you been solving this problem?

OpenStudy (anonymous):

ahh damn i used the wrong formula :/ just noticed

OpenStudy (anonymous):

hmmm? ok. What's the formula you need to use?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

ahhh I see. Well why don't you turn all the equations in the answer into the same form of \[y=mx+b\] do you know how to do so?

OpenStudy (anonymous):

I am not sure if you know this but that form is the standard form of a line. So if you have an equation of a line you should be able to turn that equation into that form.

OpenStudy (anonymous):

Well i used that formula :/ but then i noticed there aren't any y= accept in the middle

OpenStudy (anonymous):

hmmm I am a bit puzzled at what you are getting at so can you change A into y=mx+b?

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

Ahhh ok so let me walk you through it, so our equation is \[-7x+y=-44\]in order to get it to the y=mx+b form we need to move the "-7x" to the other side. To do so we can simply just add "7x" to both sides since that would cancel out the "-7x" on the left side...Does this make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if we do that what becomes of our equation?

OpenStudy (anonymous):

I hate to do this but...actually I got to go. I'll leave you with this...the equation becomes \[y=7x-44\] which is the standard form! You need to do this to all the answers to see which are different and then plug in your point in the equations and see if that point is part of the line to solve the question. If you have any questions let me know, I will answer them later.

OpenStudy (anonymous):

ah sorry , thanks !